OpenStudy (anonymous):

Use the quadratic formula to solve the equation: 3 x^2 = -3/2 x + 1

OpenStudy (ash2326):

ax^2+bx+c=0 \[x=(-b+\sqrt(b^2-4ac))/2a\] \[ x=(-b-\sqrt(b^2-4ac))/2a\] 3x^2=-3/2x+1 6x^2=-3x+2 6x^2+3x-2=0 here a =6, b=3 c=-2 so the solution can be obtained by substituting these values in the formulae so \[x=(-3+sqrt(9+4*6*2))/12\] \[x=(-3+sqrt(57))/12\] and the other solution is \[x=(-3-sqrt(57))/12\]

OpenStudy (anonymous):

What did you do to get the 6x^2?

OpenStudy (ash2326):

i multiplied the whole equation by 2